OpenStudy (anonymous):

i am just shortening the question what is the percentage of the ice cream ordered was strawberry? strawberry is 9 chocolate is 11 vanilla is 10 a. 40% b. 33 1/3% c. 30% d. 9%

OpenStudy (anonymous):

divide the number of strawberry by the total

OpenStudy (anonymous):

then turn it in to a percent by moving the decimal 2 places left

OpenStudy (anonymous):

oops make that "right"

OpenStudy (anonymous):

for example \(.27=27\%\)

OpenStudy (boldjon):

yeah i was gonna say that

OpenStudy (boldjon):

to the right i mean

OpenStudy (anonymous):

left, right, whatever

OpenStudy (boldjon):

lol

OpenStudy (anonymous):

@bebegirls \[strawberry \ is\ 9 \\chocolate \ is\ 11 \\vanilla\ is\ 10\\ total= 9+11+10=30\\ percentage\ of\ strawberry= \frac{ No.\ of\ strawberry }{ total }\times100\ \\=\frac{ 9 }{ 30} \times100\\=\frac{ 900 }{30 }\\=30 %\] Clearly option C is correct.